# Leveraging Pre-Trained Visual Transformers for Multi-Band Photometric Light Curve Classification

### v1

@@ 4 Results
Table 2: F1-score (%) on the MACHO test set for different models trained on varying sample sizes.
@@ 4.1 Model performance in single- and multi-band light curves
To evaluate the classification performance of our approach using a single band, we compared SwinV2 with Astromer. Our initial experiments focused on the R band of the MACHO dataset to ensure direct comparability with the results reported in the referenced study. We conducted hyperparameter tuning to optimize the model’s performance and to assess its ability to learn from datasets with small to medium sample sizes per class. The best combination of hyperparameters, using one band, for the 20 and 500 spc used to train the model, as well as the Full dataset, can be found at the top of Table 4 in Appendix A. These combinations were then used to generate the SwinV2 results presented in Table 2.
@@
To demonstrate the impact of the multi-band approach, we extended our analysis to incorporate both the R and B bands from the MACHO dataset, employing two distinct plotting strategies: Grid and Overlay. As before, we performed hyperparameter tuning on each scenario. The best combination of hyperparameters using two bands can be seen in the middle of Table 4 in Appendix A. The results in Fig. 5 show that incorporating multi-band inputs improves F1-scores compared to single-band experiments. Using two bands led to performance improvements of $6.8\%$, $4.3\%$, and $4.5\%$ for the 20, 500, and Full dataset settings, respectively, relative to the best multiband approach in each case. These improvements highlight the model’s ability to effectively leverage the additional information provided by multiple bands, reinforcing the value of multi-band configurations for capturing the complex behavior of light curves through image-based representations. For 20 spc, Overlay outperformed Grid (F1-score: $66.9\%$ vs. $64.3\%$). With 500 spc, Grid achieved an F1-score of $81.7\%$, marginally surpassing Overlay ($80.6\%$). When using the Full dataset, Grid attained a higher F1-score of $84.3\%$, slightly outperforming Overlay ($82.3\%$). These results indicate that Grid achieved better performance in two configurations. However, the differences were minimal.
@@ 4.2 Model performance in ELAsTiCC Challenge
To assess the model’s performance on a large-scale dataset, we used ELAsTiCC, which includes six bands. In line with prior datasets, we fine-tuned hyperparameters to enhance the model’s effectiveness. The best hyperparameter combination for Grid and Overlay can be found at the bottom of Table 4 in Appendix A.
@@ Acknowledgements.
Appendix A Hyperparameters

### v2

@@ 4 Results
Table 2: F1-score (%) on the MACHO test set for models trained with 20 and 500 samples per class (spc), and the Full dataset, comprising 20 29420\,294 labeled light curves used for training.
@@ 4.1 Model performance in single- and multi-band light curves
To evaluate the classification performance of our approach using a single band, we compared SwinV2 with Astromer. Our initial experiments focused on the R band of the MACHO dataset to ensure direct comparability with the results reported in the referenced study. We conducted hyperparameter tuning to optimize the model’s performance and to assess its ability to learn from datasets with small to medium sample sizes per class. The best combination of hyperparameters, using one band, for the 20 and 500 spc used to train the model, as well as the Full dataset, can be found at the top of Table 5 in Appendix B. These combinations were then used to generate the SwinV2 results presented in Table 2.
@@
To demonstrate the impact of the multi-band approach, we extended our analysis to incorporate both the R and B bands from the MACHO dataset, employing two distinct plotting strategies: Grid and Overlay. As before, we performed hyperparameter tuning on each scenario. The best combination of hyperparameters using two bands can be seen in the middle of Table 5 in Appendix B. The results in Fig. 5 show that incorporating multi-band inputs improves F1-scores compared to single-band experiments. Using two bands led to performance improvements of $6.8\%$, $4.3\%$, and $4.5\%$ for the 20, 500, and Full dataset settings, respectively, relative to the best multiband approach in each case. These improvements highlight the model’s ability to effectively leverage the additional information provided by multiple bands, reinforcing the value of multi-band configurations for capturing the complex behavior of light curves through image-based representations. For 20 spc, Overlay outperformed Grid (F1-score: $66.9\%$ vs. $64.3\%$). With 500 spc, Grid achieved an F1-score of $81.7\%$, marginally surpassing Overlay ($80.6\%$). When using the Full dataset, Grid attained a higher F1-score of $84.3\%$, slightly outperforming Overlay ($82.3\%$). These results indicate that Grid achieved better performance in two configurations. However, the differences were minimal.
@@
Evaluating model performance under realistic observational constraints is important to understand how models behave beyond controlled experimental settings. In our earlier experiments (see Fig. 5), we used randomly sampled subsets of the MACHO dataset with 20 and 500 samples per class to compare SwinV2 with Astromer. These results showed that SwinV2 reached higher F1-scores across different training sizes. However, those experiments did not reflect the sampling biases often found in real spectroscopic follow-up observations, where brighter sources tend to be prioritized.
Table 3: F1-score (%) on the MACHO test set using the best SwinV2 model, evaluated under brightness-biased sampling in the 20 spc-b and 500 spc-b settings. The test set remains unchanged.
Approach 20 spc-b 500 spc-b
One-band 48.9 $\pm$ 1.4 69.5 $\pm$ 1.8
Two-band (Overlay) 55.8 $\pm$ 4.9 72.9 $\pm$ 1.5
Two-band (Grid) 51.6 $\pm$ 7.0 74.7 $\pm$ 1.2
To explore this aspect, we conducted a complementary experiment using a brightness-biased sampling strategy. Specifically, we modified only the training set by selecting the brightest objects of each class from the Full dataset, based on their median magnitudes, while maintaining the 20 and 500 samples per class configurations. We refer to these settings as 20 spc-b and 500 spc-b, where the suffix -b denotes brightness-biased sampling. The test set remained unchanged from previous experiments. This introduces a distribution shift between training and testing, as the training set includes only the brightest sources, while the test set covers the full magnitude range. The results, presented in Table 3, show lower F1-scores compared to those obtained with randomly sampled training data. For 20 spc-b, the F1-score dropped from 60.1% to 48.9% in the one-band configuration, from $66.9\%$ to $55.8\%$ in the Overlay approach, and from $64.3\%$ to $51.6\%$ in the Grid approach. For 500 spc-b, the F1-score decreased from $77.4\%$ to $69.5\%$ in one-band, from $80.6\%$ to $72.9\%$ in Overlay, and from $81.7\%$ to $74.7\%$ in Grid. Despite these differences, the model still achieves relatively high scores, particularly with two-band inputs. Among the visual strategies, Overlay performs better under smaller training sizes, while Grid yields the highest performance with 500 spc-b. These findings indicate that the model remains effective even under sampling conditions that reflect realistic observational biases.
@@ 4.2 Model performance in ELAsTiCC Challenge
To assess the model’s performance on a large-scale dataset, we used ELAsTiCC, which includes six bands. In line with prior datasets, we fine-tuned hyperparameters to enhance the model’s effectiveness. The best hyperparameter combination for Grid and Overlay can be found at the bottom of Table 5 in Appendix B.
@@ Acknowledgements.
Appendix A Visual Representations of MACHO and ELAsTiCC Light Curves
Figure 10: Examples from the MACHO dataset illustrating six objects from different classes. Each object is shown in three formats: the original single-band light curve, and two-band representations generated using the Overlay and Grid approaches.
Figure 11: Examples from the ELAsTiCC dataset showing six-band light curves for six objects from different classes. Each object is represented using both the Overlay and Grid approaches.
Figures 10 and 11 present randomly selected examples from the MACHO and ELAsTiCC datasets to illustrate the visual characteristics of the light curves used in our experiments. Each figure shows six objects from different classes, represented using the approaches evaluated in our pipeline and visualized with the hyperparameters that yielded the best classification performance. For MACHO, we include three visual representations per object: the single-band light curve (One-band), and the two-band inputs constructed using the Overlay and Grid approaches. For ELAsTiCC, which contains six-band light curves, we show the corresponding Overlay and Grid representations. Within each dataset, the same objects and class labels are used across approaches to enable direct visual comparison of how different methods encode temporal and spectral information.
Appendix B Hyperparameters